OpenStudy (anonymous):

Solve for x: log base 7(x+3) = 2

jimthompson5910 (jim_thompson5910):

\[\Large \log_{7}(x+3)=2\] \[\Large x+3=7^2\] \[\Large x+3=49\] \[\Large x+3-3=49-3\] \[\Large x=46\]